# KnowEEG: Explainable Knowledge Driven EEG Classification

### v1

@@ 4.2. State-of-the-art Methods
We compare our pipeline to many state-of-the-art methods from the literature. Recent research has seen representation based deep learning methods excel on EEG data (kostas2021bendr, )(mohammadi2024eeg2rep, ). Thus we compare versus four deep learning representation based classification models. The recently developed EEG2Rep (mohammadi2024eeg2rep, ) as well as BIOT (yang2024biot, ), BENDR (kostas2021bendr, ) and MAEEG (chien2022maeeg, ). Per (mohammadi2024eeg2rep, ) these self-supervised models achieve the highest performance if they are first able to learn a representation of the EEG by training on the data in a self-supervised fashion without labels before being trained on the data with labels. This is referred to as the ’Fine Tuning’ setting. Thus, we deploy these models in this fine-tuning fashion where they perform best. We deploy EEG2Rep in both the fine-tuning fashion and the default fashion where it is deployed directly on the data without a pre-training phase. We also select the EEG Conformer (eegconformer, ) as a competitor model (deep learning model for classification). This model is adapted from transformer architecture. Finally, we select a general time series feature-based method, Catch22 (lubba2019catch22, ). These models were all implemented using publicly available code to ensure fair evaluation.
4.3. Experimental Procedure
@@
For all datasets our pipeline KnowEEG (see 3 ) and all competitor models (see 4.2 ) were trained on the training split of the dataset, had hyperparameters tuned using the validation dataset and were tested on a holdout test set. This was done for five different random seeds. Mean and standard deviation of performance metrics were used to assess model performance. For binary classification, accuracy and Area Under ROC Curve (AUROC) are selected as the performance metrics. For multi-class classification on the TUEV dataset Balanced Accuracy and Weighted F1 score are used. This is the exact same experimental procedure as in (mohammadi2024eeg2rep, ).
@@
For KnowEEG the TSFresh (christ2018time, ) package is used with ’Efficient’ settings to calculate per electrode statistics. The MNE Python package (MNE_2013, ) is used for calculating connectivity metrics. For the Fusion Forest the number of trees hyperparameter was selected from the list [50,100,200,500, 800, 1000]. Further details for the set up of KnowEEG are provided in Appendix A.
4.4. Results
@@
Table 2 shows the average performance for all models across the DREAMER, Crowdsourced, STEW, TUEV and TUAB datasets. With 10 total performance metrics across 5 datasets KnowEEG achieves best performance vs. all other models on 6/10 metrics, second best on 3/10 (with 2/10 not significantly worse than the best model) and third best on 1/10 metrics. Thus, judging only by performance, KnowEEG is the best performing model overall versus state-of-the-art competitors.
@@
Crowdsourced is the EEG task where models perform best with the task of classifying eyes open vs eyes closed EEG data. Even on this dataset, KnowEEG has the highest accuracy. For AUROC KnowEEG significantly outperforms all other models with 98.27 and performs consistently across seeds with a small standard deviation of only 0.2. Notably, the other feature based method Catch22 (lubba2019catch22, ) performs well on this dataset achieving 89.57 % accuracy and 95.92 AUROC (second best). However, Catch22 performs poorly on all other datasets illustrating that traditional feature-based methods can struggle to perform well across a diverse set of EEG classification tasks.
@@
For the TUEV dataset KnowEEG is third in balanced accuracy yet significantly outperforms all other models in weighted F1 score. TUEV is an extremely unbalanced dataset with 6 classes of EEG events. Therefore in this case balanced accuracy is not necessarily the best measure of overall model performance. This is because balanced accuracy ignores class distributions resulting in smaller classes having a disproportional impact on the balanced accuracy score and can be a drawback if targeting good accuracy on the entire dataset(grandini2020metrics, ). Weighted F1-score combines both precision and recall with each class weighted proportionally. Thus, it is notable that KnowEEG outperforms all other models in this metric significantly.
For abnormal EEG classification on TUAB KnowEEG performs second best to EEG2Rep on both accuracy and AUROC. However KnowEEG is within standard deviation of EEG2Rep so this result is not significant. BIOT also reaches the same performances as EEG2Rep and KnowEEG when accounting for standard deviation across runs.
EEG2Rep Pre-trained (mohammadi2024eeg2rep, ) is overall the second best performing model achieving best performance or second best performance on 8/10 metrics across the five datasets. Besides performance, when compared with Deep Learning models, KnowEEG has the significant advantage of not requiring GPU for training. All self-supervised representation-based models (except EEG2Rep Default) were trained first on the training data in self-supervised fashion and then tuned in supervised fashion. This requires significant GPU resource, in particular for the larger datasets. KnowEEG does not require any GPU resource. KnowEEG also has the significant advantage of using calculated features that are inherently interpretable and a tree-based model which enables users to access feature importances directly. We demonstrate the benefits of this explainability in Section 4.5.
Table 2. Performance of KnowEEG versus competitor models across datasets. KnowEEG results are shaded in blue. Best performance per metric per dataset is in bold with second best underlined.
@@
Acc AUROC Acc AUROC Acc AUROC B-Acc W-F1 Acc AUROC
BENDR 54.45$\pm$2.11 53.02$\pm$1.31 83.78$\pm$2.35 83.80$\pm$2.63 69.74$\pm$2.11 69.77$\pm$2.03 41.17$\pm$2.89 67.31$\pm$2.96 76.96$\pm$3.98 83.97$\pm$3.44
MAEEG 53.63$\pm$2.61 52.08$\pm$2.36 86.75$\pm$3.50 86.21$\pm$3.41 72.46$\pm$3.67 72.50$\pm$3.22 41.23$\pm$3.65 67.38$\pm$3.69 77.56$\pm$3.56 86.56$\pm$3.33
BIOT 53.45$\pm$2.01 53.53$\pm$1.82 87.95$\pm$3.52 87.78$\pm$3.09 69.88$\pm$2.15 70.11$\pm$2.57 46.02$\pm$1.68 69.98$\pm$1.99 79.21$\pm$2.15 87.42$\pm$2.01
Catch22 59.06 $\pm$1.61 51.6$\pm$2.17 89.57 $\pm$0.64 95.82 $\pm$0.3 55.72 $\pm$0.2 53.2$\pm$0.14 33.65 $\pm$0.5 67.01 $\pm$0.34 69.84 $\pm$1.06 66.16 $\pm$1.12
EEGConformer 55.40 $\pm$2.76 54.19 $\pm$2.57 87.62 $\pm$2.91 87.24 $\pm$2.06 74.35$\pm$1.88 73.11 $\pm$1.95 43.88$\pm$1.62 67.20$\pm$2.31 78.36 $\pm$1.71 83.89 $\pm$2.64
EEG2Rep (Default) 54.61$\pm$2.22 53.61$\pm$2.07 91.19$\pm$1.18 91.22$\pm$1.23 70.26$\pm$1.59 69.77$\pm$2.03 44.25$\pm$3.01 68.95$\pm$2.89 77.85$\pm$3.14 84.91$\pm$3.07
EEG2Rep (Pre-Trained) 60.37$\pm$1.52 59.42$\pm$1.45 94.13$\pm$1.21 94.13$\pm$2.17 73.60$\pm$1.47 74.40$\pm$1.50 52.95$\pm$1.58 75.08$\pm$1.21 80.52$\pm$2.22 88.43$\pm$3.09
KnowEEG 61.53$\pm$2.59 55.29$\pm$2.58 94.53$\pm$0.35 98.27$\pm$0.13 77.86$\pm$0.19 85.96$\pm$0.8 45.06$\pm$0.71 78.28$\pm$0.56 80.18$\pm$0.24 87.71$\pm$0.23
4.5. KnowEEG Explainability Analysis
Here, we analyze the KnowEEG model trained on the training and validation sets for Crowdsource and demonstrate the key explainability benefits. The Crowdsource dataset presents the problem of binary classification on 14-channel EEG data. The two classes are eyes open and eyes closed. For interpretability, we propose analyzing the connectivity and statistical features separately.
In hyperparameter selection, Coherence was selected as the connectivity measure for the Crowdsource dataset. As described in Section 3, the selected connectivity metric is calculated between each electrode pair over six defined frequency bands (alpha, theta, delta, sigma, beta, gamma). Therefore, there are 91 x 6 = 548 coherence features. The number 91 refers to every combination of electrode-electrode pairs for the 14 electrodes (excluding pairing an electrode with itself).
A simple first step in analyzing these feature importances is to plot their distribution. Figure 2 shows an exponential like feature distribution with many feature importances with low values close to zero and fewer and fewer features with higher importances.
@@ 4.5. KnowEEG Explainability Analysis
Histogram plot showing distribution of feature importances Refer to caption
Figure 2. Feature importances histogram for connectivity features (Coherence) on Crowdsource dataset.
We propose as a simple next step to analyze connectivity feature importances per power band. Due to interactions between features, summing feature importances per band does not accurately represent group feature importance ((NEURIPS2020_c7bf0b7c, )). Therefore, instead we compare each channel-channel connectivity measure per band and rank them. For example, electrodes AF3-AF4 will have 6 coherence connectivity values, one for each band. The band with the highest feature importance for each electrode-electrode pair receives a score of 5 down to 6th place which receives a score of 0. The per band ranking feature importance scores are shown in figure 3.
@@
Bar chart showing the per power band feature importance score Refer to caption
Figure 3. Feature Importance score per power band for coherence connectivity measure on the Crowdsource dataset.
Per figure 3, the alpha band has the highest connectivity importance score followed by gamma and beta with delta, theta and sigma having the lowest importance rank scores. This suggests that coherence in the alpha band is most discriminative versus other bands for eyes closed versus eyes open classification. We can take this analysis further by visualizing alpha coherence for the two classes. We can do this on a surface plot of the 14 electrodes on the head, where each electrode is plotted as the sum of alpha connectivity values to all other electrodes.
@@
Heatmap of the head showing alpha coherence for eyes closed and eyes open. Refer to caption
Figure 4. Alpha coherence across participants in training and validation set for eyes closed and eyes open plotted as heatmap on the head.
In figure 4 we plot the mean alpha coherence for the eyes open and eyes closed classes. The dark colours for eyes open show low coherence across the entire head for alpha coherence in contrast to higher alpha coherence for the entire head for eyes closed participants. Alpha coherence is particularly high for the frontal region of the brain for eyes closed participants. This result is consistent with existing EEG literature which shows that ”functional connectivity in the alpha band decreases in the eyes open condition compared to eyes closed” (gomezram_eyes_op_close, ). This is referred to as alpha de-synchronization and demonstrates the effectiveness of our pipeline. Without sacrificing performance, we are able to explain how our model arrives at results and correctly discover knowledge about our classes. As implied by the feature importances graph, the corresponding heat map plots for the other power bands show noticeable differences between the two classes for gamma and beta and smaller differences for delta, theta and sigma. For reference the lowest importance band plot (theta) is provided in Appendix D.
We must note that in Figure 3 there are 91 features (electrode-electrode coherence in the alpha band) summarized on a 14 electrode plot of the head. Therefore, there will not necessarily always be a clear distinction between the classes when generating a summary plot in this fashion and one may need to generate more specific plots. However, our KnowEEG pipeline will always show the raw feature importances and enable you to generate visually discriminative plots for your classes when the classifier (KnowEEG) has been able to differentiate the classes.
In a similar fashion to the per band analysis for connectivity one could also do a per channel-channel analysis to determine if the connectivity between any channel pairs is particularly discriminative for the classification problem. In doing this, we do not discover any noteworthy patterns and for conciseness proceed to analyzing the statistical features.
The feature importances for per channel statistics also follow an exponential-like distribution with many features having small importances close to 0 and fewer features having higher importances. As a simple first step we can analyze the top n features by feature importance. We select n = 10 features presenting these features in Table 3 below.
@@
Table showing the 10 highest importance features Rank Feature Parameters Channel Brain Region 1 Permutation Entropy Dimension 6 Tau 1 O1 Occipital 2 Permutation Entropy Dimension 4 Tau 1 O2 Occipital 3 Permutation Entropy Dimension 5 Tau 1 O2 Occipital 4 Partial Autocorrelation Lag 2 O1 Occipital 5 Number Peaks Support 1 O2 Occipital 6 Permutation Entropy Dimension 5 Tau 1 O1 Occipital 7 Permutation Entropy Dimension 3 Tau 1 O2 Occipital 8 Mean Absolute Change N/A FC6 Right Central 9 Permutation Entropy Dimension 7 Tau 1 O1 Occipital 10 Quantile Change Mean Upper 0.8, Lower 0 FC6 Right Central
@@
Table 3. Statistical features with highest importance ranked 1 to 10. Table displays feature name, parameters specifying feature calculation, corresponding electrode and corresponding brain region.
@@
From Table 3 it is notable that 7 out of the top 7 highest importance features are from the Occipital brain region. This suggests that the Occipital brain region provides features that are most discriminative of the two classes. The two classes are eyes open and eyes closed. We know from existing neuroscience research that the Occipital region of the brain is responsible for visual processing as it ”houses the visual cortex responsible for processing and interpreting visual stimuli” (fossa2024occipital, ). Therefore, again, from our KnowEEG pipeline we have been able to discover knowledge (signals from the Occipital brain region being most discriminative for the two classes) and have verified this versus existing neuroscience literature. This shows that our pipeline can aid discovery of new information where differences between classes are not yet understood.
@@
[…]
Bar chart showing count of feature importances per Brain Region Refer to caption
Figure 5. Count of features in top N importances (N = 5, 10, 20, 100) by Brain Region
Another interesting finding from Table 3 is that the three highest importance features and 6 of the top 10 are permutation entropy from electrodes in the Occipital brain region. Permutation entropy is considered as ’a natural complexity measure for time series’ (bandt2002permutation, ). The calculation of permutation entropy requires two parameters (Dimension and Tau) which is why this feature can appear multiple times for the same electrode in Table 3.
We plot the Kernel Density Estimates for eyes closed and eyes open for the highest importance ranked feature in Figure 6. It is clear from the figures that the eyes closed group has a much broader distribution compared with the eyes open group. The eyes open group has a narrower (standard deviation 0.16 vs 0.47) and slightly higher peak. When testing the distributions for this feature, we use the two sample Kolmogorov–Smirnov test we find that the two distributions vary significantly (p-value<0.0001).
[…]
KDE plot showing distribution of permutation entropy feature on O1 Channel for eyes open and eyes closed classes Refer to caption
Figure 6. KDE for Permutation Entropy of EEG from O1 Channel (Occipital Lobe) for eyes open and eyes closed classes. Corresponds to rank 1 feature from Table 3

### v2

@@ 4.2 State-of-the-art Methods
We compare our pipeline to many state-of-the-art methods from the literature. Recent research has seen representation based deep learning methods excel on EEG data [25][32]. Thus we compare versus four deep learning representation based classification models. The recently developed EEG2Rep [32] as well as BIOT [42], BENDR [25] and MAEEG [10]. Per [32] these self-supervised models achieve the highest performance if they are first able to learn a representation of the EEG by training on the data in a self-supervised fashion without labels before being trained on the data with labels. This is referred to as the ‘Fine Tuning’ setting. Thus, we deploy these models in this fine-tuning fashion where they perform best. We deploy EEG2Rep in both the fine-tuning fashion and the default fashion where it is deployed directly on the data without a pre-training phase. We also select the EEG Conformer [38] as a competitor model (deep learning model for classification). This model is adapted from transformer architecture. Finally, we select a general time series feature-based method, Catch22 [30]. These models were all implemented using publicly available code to ensure fair evaluation.
4.3 Experimental Procedure
@@
For all datasets our pipeline KnowEEG (see 3 ) and all competitor models (see 4.2) were trained on the training split of the dataset, had hyperparameters tuned using the validation dataset and were tested on a holdout test set. This was done for five different random seeds. Mean and standard deviation of performance metrics were used to assess model performance. For binary classification, accuracy and Area Under ROC Curve (AUROC) are selected as the performance metrics. For multi-class classification on the TUEV dataset Balanced Accuracy and Weighted F1 score are used. This is the exact same experimental procedure as in [32].
@@
For KnowEEG the TSFresh [12] package is used with ‘Efficient’ settings to calculate per electrode statistics. The MNE Python package [18] is used for calculating connectivity metrics. For the Fusion Forest the number of trees hyperparameter was selected from the list [50,100,200,500, 800, 1000]. Further details for the set up of KnowEEG are provided in A.
4.4 Results
@@
Table 2 shows the average performance for all models across the DREAMER, Crowdsourced, STEW, TUEV and TUAB datasets. With 10 total performance metrics across 5 datasets KnowEEG achieves best performance vs. all other models on 5/10 metrics, second best on 4/10 (with 2/10 not significantly worse than the best model) and third best on 1/10 metrics. Thus, judging by performance ranking alone, KnowEEG is the best performing model overall versus state-of-the-art competitors.
@@
Crowdsourced is the EEG task where models perform best with the task of classifying eyes open vs eyes closed EEG data. On this dataset, KnowEEG has the second highest accuracy with 92.81 % performing marginally worse than the pre-trained EEG2Rep model. For AUROC KnowEEG outperforms all other models with 96.72 and performs consistently across seeds with a small standard deviation of only 0.21. Notably, the other feature based method Catch22 [30] performs well on this dataset achieving 89.57 % accuracy and 95.82 AUROC (second best). However, Catch22 performs poorly on all other datasets illustrating that traditional feature-based methods can struggle to perform well across a diverse set of EEG classification tasks.
@@
For the TUEV dataset KnowEEG is third in balanced accuracy yet significantly outperforms all other models in weighted F1 score. TUEV is an extremely unbalanced dataset with 6 classes of EEG events. Therefore in this case balanced accuracy is not necessarily the best measure of overall model performance. This is because balanced accuracy ignores class distributions resulting in smaller classes having a disproportional impact on the balanced accuracy score and can be a drawback if targeting good accuracy on the entire dataset[19]. Weighted F1-score combines both precision and recall with each class weighted proportionally. Thus, it is notable that KnowEEG outperforms all other models in this metric significantly.
For abnormal EEG classification on TUAB KnowEEG performs second best to EEG2Rep on both accuracy and AUROC. However KnowEEG is within the standard deviation of EEG2Rep so this result is not significant. BIOT also reaches the same performances as EEG2Rep and KnowEEG when accounting for standard deviation across runs.
EEG2Rep Pre-trained [32] is overall the second best performing model achieving best performance or second best performance on 8/10 metrics across the five datasets. Besides performance, when compared with Deep Learning models, KnowEEG has the significant advantage of not requiring GPU for training. All self-supervised representation-based models (except EEG2Rep Default) were trained first on the training data in self-supervised fashion and then tuned in supervised fashion. This requires significant GPU resource, in particular for the larger datasets. KnowEEG does not require any GPU resource. KnowEEG also has the significant advantage of using calculated features that are inherently interpretable and a tree-based model which enables users to access feature importances directly. We demonstrate the benefits of this explainability in Section 4.5.
Table 2: Performance of KnowEEG versus competitor models across datasets. KnowEEG results are shaded in blue. Best performance per dataset is in bold with second best underlined with Accuracy results in table a) and AUROC in table b).
(a) Accuracy (Acc) Performance
@@
Acc Acc Acc B-Acc Acc
@@
[…]
(b) AUROC Performance
Models DREAMER Crowdsourced STEW TUEV TUAB
AUROC AUROC AUROC W-F1 AUROC
[…]
4.5 KnowEEG Explainability Analysis
Here, we analyse the KnowEEG model trained on the training and validation sets for the Crowdsource dataset and demonstrate the key explainability benefits. The Crowdsource task presents the problem of binary classification on 14-channel EEG data. The two classes are eyes open and eyes closed. For interpretability, we propose analysing the connectivity and statistical features separately.
In hyperparameter selection, Functional Power Connectivity (FPC) was selected as the connectivity measure for the Crowdsource dataset. As defined in B, the FPC feature set is calculated between each electrode pair over six defined frequency bands (alpha, theta, delta, sigma, beta, gamma) as well as having additional features indicating relative band power values per electrode. The FPC features therefore describe both band power activity and functional connectivity across the brain.
A first step in analysing these feature importances is to plot their distribution. Figure 2 shows an exponential like feature distribution with many feature importances with low values close to zero and fewer and fewer features with higher importances.
Figure 2: This histogram displays the Gini importance values for the Functional Power Connectivity features (FPC) used by the trained KnowEEG model on the Crowdsource dataset. The x-axis represents the feature importance score with the y-axis showing the count of features at that importance level. The distribution is highly skewed with an exponential-like decay, showing that a large majority of features have an importance score near zero, while a small group of features have higher importance scores.
We propose as a simple next step to analyse connectivity feature importances per power band. Due to interactions between features, summing feature importances per band does not accurately represent group feature importance [13]. Therefore, instead we compare each channel-channel connectivity measure per band and rank them. For example, electrodes AF3-AF4 will have 6 FPC connectivity values, one for each band. The band with the highest feature importance for each electrode-electrode pair receives a score of 5 down to 6th place which receives a score of 0. The per band ranking feature importance scores are shown in figure 3.
Figure 3: Feature Importance score per power band for Functional Power Connectivity on the Crowdsource dataset. FPC gamma features have the highest feature importance rank score overall followed by alpha, delta, theta and sigma with the lowest score.
Per figure 3, the Functional Power Connectivity features on the gamma band have the highest importance followed by alpha and then delta, beta, theta and sigma with the lowest importance rank. This suggests that FPC features in gamma and alpha bands are most discriminative versus other bands for eyes closed versus eyes open classification. We can take this analysis further by visualizing FPC features for alpha and gamma for the two classes. We can do this on a surface plot of the 14 electrodes on the head, with the FPC band power features separated from FPC pure connectivity features. We plot FPC band power in figure 4 and connectivity in figure 5.
Figure 4: This plot shows mean relative band power in the alpha and gamma bands for eyes closed (left) and eyes open (right) participants. For the Alpha band the eyes closed group have a higher mean power particular in the Occipital Lobes towards the back of the head. For the Gamma band the eyes open group have a higher mean power with peaks in the Left and Right Temporal / Central regions of the brain.
Figure 5: This plot shows the mean Alpha and Gamma band connectivity for the eyes closed (left) and eyes open (right) groups per the FPC features (defined in B). The eyes closed group have higher Alpha connectivity across the brain particularly in the Occipital regions. For Gamma connectivity the differences between the groups are not visually significant.
Per figure 4 the mean eyes open alpha band power is noticeably higher in the eyes closed group vs. eyes open, particularly in the Occipital electrodes towards the back of the head. This result is in line with the neuroscience literature that show an increase in alpha power in the eyes closed state [17, 4]. Contrastively the mean gamma band power is significantly higher across the head for the eyes open state vs. eyes closed per figure 4. Again, this is an expected result supported by EEG studies [16]. The eyes being open increases alertness and raises arousal in the brain resulting in stronger activity in the gamma band.
Figure 5 shows that mean FPC alpha connectivity is higher for the eyes closed participants vs. eyes open. This phenomenon is known as alpha de-synchronization and is consistent with EEG literature that shows ”functional connectivity in the alpha band decreases in the eyes open condition compared to eyes closed” [17]. Mean FPC gamma connectivity does not show significant differences between the eyes open and closed groups. This likely reflects the low amplitude and high variability of resting-state gamma activity, making gamma band connectivity a weak discriminator for distinguishing between the eyes open and closed states.
In summary, from our analysis of the Functional Power Connectivity features we have uncovered known neuroscience knowledge. Specifically, we have found increased mean alpha band activity and connectivity in the eyes closed state compared to eyes open as well as increased mean gamma band activity in the eyes open class [17, 23, 16, 4]. This ability of our pipeline to uncover knowledge about the classes is a significant strength. We must note though, that the summary plots showing mean values across the head for both classes will not always show a clear distinction between classes. However, they are a useful summary view of the features and the detailed KnowEEG feature importance scores will always be available to guide more detailed analysis of the classes.
Furthermore, a granular per channel analysis could also be completed to determine if the connectivity between any channel pairs is particularly discriminative for the classification problem. To keep the analysis focused we move on to the per channel statistical features.
The feature importances for per channel statistics also follow an exponential-like distribution with many features having small importances close to 0 and fewer features having higher importances. As a simple first step we can analyse the top n features by feature importance. We select n = 10 features presenting these features in Table 3 below.
@@ 4.5 KnowEEG Explainability Analysis
Table 3: Statistical features with highest importance ranked 1 to 10. The table displays the feature name, parameters specifying feature calculation, corresponding electrode and corresponding brain region. It is notable that the top 7 features are all from the Occipital brain region.
Rank Feature Parameters Channel Brain Region
1 Permutation Entropy Dimension 6 Tau 1 O1 Occipital
2 Permutation Entropy Dimension 4 Tau 1 O2 Occipital
3 Permutation Entropy Dimension 5 Tau 1 O2 Occipital
4 Partial Autocorrelation Lag 2 O1 Occipital
5 Number Peaks Support 1 O2 Occipital
6 Permutation Entropy Dimension 5 Tau 1 O1 Occipital
7 Permutation Entropy Dimension 3 Tau 1 O2 Occipital
8 Mean Absolute Change N/A FC6 Right Central
9 Permutation Entropy Dimension 7 Tau 1 O1 Occipital
10 Quantile Change Mean Upper 0.8, Lower 0 FC6 Right Central
@@ 9 Permutation Entropy Dimension 7 Tau 1 O1 Occipital
From Table 3 it is notable that 7 out of the top 7 highest importance features are from the Occipital brain region. This suggests that the Occipital brain region provides features that are most discriminative of the two classes. The two classes are eyes open and eyes closed. We know from existing neuroscience research that the Occipital region of the brain is responsible for visual processing as it ”houses the visual cortex responsible for processing and interpreting visual stimuli” [1]. Therefore, again, from our KnowEEG pipeline we have been able to discover knowledge (signals from the Occipital brain region being most discriminative for the two classes) and have verified this versus existing neuroscience literature. This shows that our pipeline can aid discovery of new information where differences between classes are not yet understood.
@@
Figure 6: Count of features in top N importances (N = 5, 10, 20, 100) by Brain Region. From this bar chart it is clear that the Occipital region has a disproportionately large number of highly discriminative features vs. other regions. The chart shows that 5 out of the top 5 (dark blue) most important features are from the Occipital region as well as 7 of the top 10, 13 of the top 20 and 32 of the top 100.
Another interesting finding from Table 3 is that the three highest importance features and 6 of the top 10 are permutation entropy from electrodes in the Occipital brain region. Permutation entropy is considered as ‘a natural complexity measure for time series’ [3]. The calculation of permutation entropy requires two parameters (Dimension and Tau) which is why this feature can appear multiple times for the same electrode in Table 3.
We plot the Kernel Density Estimates for eyes closed and eyes open for the highest importance ranked feature in Figure 7. It is clear from the figures that the eyes closed group has a much broader distribution compared with the eyes open group. The eyes open group distribution is narrower (standard deviation 0.16 vs 0.47) with a slightly higher peak. When testing the distributions for this feature, we use the two sample Kolmogorov–Smirnov test and find that the two distributions vary significantly (p-value<0.0001).
Figure 7: KDE plot of the rank 1 feature from Table 3: Permutation Entropy of the O1 (Occipital Lobe) EEG channel. The distribution for the ‘eyes open’ class (green) is distinctly narrower and has a higher peak density, while the ‘eyes closed’ class (blue) shows a broader distribution with a wider spread of values.